Suppose market demand is given Qd=a-bp as and market supply is given as, Qs=c+dp and then find the following



o Equilibrium price



o Equilibrium quantity



o Price elasticity of demand at equilibrium



o Price elasticity of supply at equilibrium

1
Expert's answer
2021-11-24T19:28:35-0500

At equilibrium price, the quantity demanded is equal to quantity supplied in the market

"Q{d}=Q{s}"

therefore

"(a- bp)= (c+dp)"

"(a-c) =(dp + bp)"

"(a - c)=p(d +p)"

"p=(a - c)\/(b + d)"

"equilibrium price =(a - c)\/(b + d)"


Equilibrium quantity is gotten by substituting the equilibrium price into the demand equation

"Qd=a - bp"

"qd =a- b[\frac{ a-c}{b+d}]"

"q=a- [\frac{ab +cb}{b+d}]"


"Q =\frac{(ab+bd)-(ab+cb)}{b+d}"


"Q=\frac{ab -ab+cb+bd}{b+d}"

EQUILLIBRIUM QUANTITY IS GIVEN BY

"Q=\frac{cb + ad}{b + d}"


Price elasticity of demand at equilibrium


"price elasticity=\frac{percentage change in quantity}{percentagechange in price}"


PERCENTAGE CHANGE IN QUANTINTY ="\frac {\Delta Q}{Q}"

PERCENTAGE CHANGE IN PRICE="\frac {\Delta P}{P}"

change in quantity"\Delta Q =Q{1}- Q"

change in price"\Delta Q =P{1}- P"


PRICE ELASTICITY ="\frac{\Delta Q}{ Q} \times \frac{P}{\Delta P}"


PRICE ELASTICITY ="\frac{\Delta Q}{\Delta P } \times \frac{P}{Q}"

but from the demand equation "\frac{\Delta Q}{\Delta P }" is the gradient of the equation

therefore from "Q{s}=(a- bp)" "-b" is the gradient of the equation



the price elasticity of demand will be


"P{e}=-b \times \frac{P}{q}"